OpenStudy (anonymous):

Is it possible to do x^2 + x?

Miracrown (miracrown):

do in what sense?

OpenStudy (anonymous):

I'm doing an equation to figure out what x equals and I've backtracked to the point of \[x ^{2}+x =\frac{ 2 }{ 27 }\] but yeah, I'm not sure what to do from there...

Miracrown (miracrown):

get rid of the fraction, bring it over to LHS, see if it neatly factors, if not apply quadratic formula

OpenStudy (anonymous):

Ohhh okay thanks :)

Miracrown (miracrown):

Quadratic formula tell us that x = (-b +- Sqrt[b^2 - 4ac])/(2a) for our equation 27x^2+27x-2 = 0. a = 27, b = 27, c = -2. Therefore, x = -27 +- Sqrt[27^2 - 4 * 27 * (-2)] all over (2*27) Divide out the 27s so -27/(2*27) = -1/2 And we can bring the 1/27 inside the square root by making is Sqrt[1/27^2] Therefore, x = -1/2 +- 1/2 Sqrt[ (27^2+4*(27)*2) / 27^2] where we've canceled the minus signs of the 4 and 2 originally to make a + sign Recognize that there's a 27 in both terms in the numerator (the original reason I did it this way) the result is x = -1/2 +- 1/2 Sqrt[1+8/27] Simplify the fraction, rationalize it and you get the final result x= 1/2(-1+- Sqrt[105]/9) clear? @freya.13

OpenStudy (anonymous):

Thank you so much ! :D haha

Miracrown (miracrown):

You're welcome. :)

OpenStudy (anonymous):

Lol what grade is quadratic formula usually learned in cos I think i shld know it but i dont OoO

Miracrown (miracrown):

hm..I first learned this in 7th grade, but I do not know if that has an universal applicability or applicability to present times. In the past 10 years these sorts of curriculums have changed a lot, I believe.

Miracrown (miracrown):

I know there are foreign students who learn this in grade school foreign in the sense of not from the USA originally

Miracrown (miracrown):

I think it should be taught in grade school because I think kids can handle it ! :)

OpenStudy (anonymous):

haha well I'm in Australia and grade 9... lol oh well o.o